OpenStudy (anonymous):

Find the slope of the line passing through the points (-3, 4) and (2, -1).

OpenStudy (chriss):

\[\frac{(-1)-4}{2-(-3)}\] \[\frac{-5}{5}\] -1

OpenStudy (anonymous):

it is the y coordinate over the x coordinate (4-(-1)/(-3)-2 =-1